The surface area of a certain cube is
600 square inches.
What is the new surface area of the cube if
the length of each edge is reduced by half

Answer by jim_thompson5910(35256)

SA = 6s^2


600 = 6s^2


600/6 = s^2


100 = s^2


s^2 = 100


s = sqrt(100)


s = 10


So the old cube has side lengths of 10 inches. Now reduce each edge by half to get 10*(1/2) = 5. The new cube has side lengths of 5 inches.


Now compute the surface area of the smaller cube.


SA = 6s^2


SA = 6*5^2


SA = 6*25


SA = 150


The new surface area is 150 square inches. Notice how this is 1/4 of the original surface area and how 1/4 = (1/2)^2